I need to know the slope and y intercept of 2y-x=4 and how to get it. I do not grasp this concept and desperately need to know it.

Answer by Alan3354(69443)

I need to know the slope and y intercept of 2y-x=4 and how to get it. I do not grasp this concept and desperately need to know it.
================
You solve for y.
2y-x=4
2y = x + 4
y = (1/2)x + 2
Now it's in slope-intercept form, y = mx + b
---
The slope m is the coefficient of x: m = 1/2
The y-intercept is 2. b = 2 or (0,2)
That's all.